solve using the elimination method.
4a+6b=-1
a-3b=2

Answer by london maths tutor(243)

4a+6b=-1 ... equation 1
a-3b=2 ... equation 2
equation 2 x 4 :
a(4) - 3b(4) = 2(4)
4a -12b = 8
Equation 1 - Equation 2
4a + 6b - [4a - 12 b] = -1 - 8
6b + 12b = -9
18b=-9
b = -9/18
b = -1/2
Put b=-1/2 into a-3b = 2
a - 3(-1/2) = 2
a -(-3/2) = 2
a + 3/2 = 2
a = 2 - 3/2
a = 2 - 1 1/2
a = 1/2
Answer : a = 1/2 and b = -1/2